# Lightweight and Interpretable Transformer via Mixed Graph Algorithm Unrolling for Traffic Forecast

### v2

@@ Abstract
Unlike conventional “black-box” transformers with classical self-attention mechanism, we build a lightweight and interpretable transformer-like neural net by unrolling a mixed-graph-based optimization algorithm to forecast traffic with spatial and temporal dimensions. We construct two graphs: an undirected graph ${\mathcal{G}}^{u}$ capturing spatial correlations across geography, and a directed graph ${\mathcal{G}}^{d}$ capturing sequential relationships over time. We predict future samples of signal $\mathbf{x}$}, assuming it is “smooth” with respect to both ${\mathcal{G}}^{u}$ and ${\mathcal{G}}^{d}$, where we design new $\ell_{2}$ and $\ell_{1}$-norm variational terms to quantify and promote signal smoothness (low-frequency reconstruction) on a directed graph. We design an iterative algorithm based on alternating direction method of multipliers (ADMM), and unroll it into a feed-forward network for data-driven parameter learning. We insert graph learning modules for ${\mathcal{G}}^{u}$ and ${\mathcal{G}}^{d}$ that play the role of self-attention. Experiments show that our unrolled networks achieve competitive traffic forecast performance as state-of-the-art prediction schemes, while reducing parameter counts drastically. Our code is available in https://github.com/SingularityUndefined/Unrolling-GSP-STForecast.
@@ 1 Introduction
Transformer, based on the classical self-attention mechanism (Vaswani et al., 2017), is now a dominant deep learning (DL) architecture that has achieved state-of-the-art (SOTA) performance across multiple fields (Woźniak et al., 2020; Zamir et al., 2022). However, like other “off-the-shelf” DL models, transformer has a massive number of parameters, and its internal structure is not easily interpretable. Parameter reduction is a major industrial concern for practical computation and/or memory-constrained environments [^1] See a brief review of lightweight learning models for different applications in Appendix A.1.. Algorithm unrolling (Monga et al., 2021) offers an alternative hybrid model-based / data-driven paradigm; by “unrolling” iterations of a model-based algorithm minimizing a well-defined optimization objective into neural layers stacked together to form a feed-forward network for data-driven parameter learning, the resulting network can be both mathematically interpretable [^2] Common in algorithm unrolling (Monga et al., 2021), “interpretability” here means that each neural layer corresponds to an iteration of an optimization algorithm minimizing a mathematically-defined objective. and competitive in performance. Notably, Yu et al. (2023) designs an algorithm minimizing a sparse rate reduction (SRR) objective that unrolls into a transformer-like neural net for image classification. Thuc et al. (2024) designs graph-based algorithms minimizing graph smoothness priors that unroll into transformer-like neural nets for image interpolation while reducing parameters. However, only a positive undirected graph model was used in Thuc et al. (2024) to capture simple pairwise correlations between neighboring pixels in a static image.
@@
Among model-based methods, one approach to traffic forecast is to filter the signal along the spatial and temporal dimensions independently (Ramakrishna et al., 2020)e.g., employ graph spectral filters in the graph signal processing (GSP) literature (Ortega et al., 2018; Cheung et al., 2018) along the irregular spatial dimension, and traditional Fourier filters along the regular time dimension. Another approach is to model the spatial correlations across time as a sequence of slowly time-varying graphs, and then filter each spatial signal at time $t$ separately, where the changes in consecutive graph adjacency matrices in time are constrained by the Frobenius norm (Kalofolias et al., 2017), $\ell_{1}$-norm (Yamada et al., 2019), or low-rankness (Bagheri et al., 2024). However, neither approach exploits spatial and temporal relationships simultaneously for optimal performance.
@@
Recent DL-based efforts address traffic forecast by building elaborate transformer architectures to capture spatial and temporal relations (Feng & Tassiulas, 2022; Gao et al., 2022; Jin et al., 2024) [^3] A detailed review of DL models on traffic forecast is included in Appendix A.2.. Like transformers used in other fields (Zamir et al., 2022), these are also black boxes with large parameter counts. Graph attention networks (GAT) (Velickovic et al., 2018) and graph transformers (Dwivedi & Bresson, 2021) are adaptations of the self-attention mechanism in transformers for graph-structured data (e.g., computing output embeddings using only input embeddings in local graph neighborhoods). However, they are still uninterpretable and maintain large parameter sizes.
@@
In contrast, we extend Thuc et al. (2024) to build lightweight “white-box” transformers that are fully interpretable via mixed-graph algorithm unrolling for traffic forecast. Specifically, we learn two graphs from data: i) undirected graph ${\mathcal{G}}^{u}$ to capture spatial correlations across geography, and ii) a directed graph ${\mathcal{G}}^{d}$ to capture sequential relationships over time. We show that our graph learning modules play the role of self-attention (Bahdanau et al., 2014), and thus our unrolled graph-based neural nets are transformers. Given the two learned graphs, we define a prediction objective for future samples in signal $\mathbf{x}$}, assuming $\mathbf{x}$} is “smooth” with respect to (w.r.t.) both ${\mathcal{G}}^{u}$ and ${\mathcal{G}}^{d}$ in variational terms: graph Laplacian regularizer (GLR) (Pang & Cheung, 2017) for undirected ${\mathcal{G}}^{u}$, and newly designed directed graph Laplacian regularizer (DGLR) and directed graph total variation (DGTV) for directed ${\mathcal{G}}^{d}$. We design a corresponding linear-time optimization algorithm based on alternating direction method of multipliers (ADMM) (Boyd et al., 2011) that unrolls into neural layers for parameter learning. For the first time, our DGLR and DGTV variational terms quantify and promote signal smoothness on a directed graph, with intuitive low-pass filter interpretations. Experiments show that our networks achieve competitive traffic forecast performance as SOTA prediction schemes, while employing drastically fewer parameters (6.4% of transformer-based PDFormer (Jiang et al., 2023)).
@@ 2 Preliminaries
Denote by ${\mathcal{G}}^{u}({\mathcal{V}},{\mathcal{E}}^{u},{\mathbf{W}}^{u})$ an undirected graph with node set ${\mathcal{V}}=\{1,\ldots,N\}$ and undirected edge set ${\mathcal{E}}^{u}$, where $(i,j)\in{\mathcal{E}}^{u}$ implies that an undirected edge exists connecting nodes $i$ and $j$ with weight $w_{i,j}^{u}=W_{i,j}^{u}$, and $(i,j)\notin{\mathcal{E}}$ implies $W_{i,j}^{u}=0$. ${\mathbf{W}}^{u}\in\mathbb{R}^{N\times N}$ is the symmetric adjacency matrix for ${\mathcal{G}}^{u}$. Denote by ${\mathbf{D}}^{u}\in\mathbb{R}^{N\times N}$ a diagonal degree matrix, where $D_{i,i}^{u}=\sum_{j}W_{j,i}^{u}$. The symmetric graph Laplacian matrix is ${\mathbf{L}}^{u}\triangleq{\mathbf{D}}^{u}-{\mathbf{W}}^{u}$ (Ortega et al., 2018). ${\mathbf{L}}^{u}$ is positive semi-definite (PSD) if all edge weights are non-negative $w^{u}_{i,j}\geq 0,\forall i,j$ (Cheung et al., 2018). The symmetric normalized graph Laplacian matrix is ${\mathbf{L}}^{u}_{n}\triangleq{\mathbf{I}}-({\mathbf{D}}^{u})^{-1/2}{\mathbf{W}}^{u}({\mathbf{D}}^{u})^{-1/2}$, while the asymmetric random-walk graph Laplacian matrix is ${\mathbf{L}}^{u}_{r}\triangleq{\mathbf{I}}-({\mathbf{D}}^{u})^{-1}{\mathbf{W}}^{u}$.
One can define a graph spectrum by eigen-decomposing the real and symmetric graph Laplacian ${\mathbf{L}}^{u}$ (or normalized graph Laplacian ${\mathbf{L}}_{n}^{u}$), where the $k$-th eigen-pair $(\lambda_{k},{\mathbf{v}}_{k})$ is interpreted as the $k$-th graph frequency and Fourier mode, respectively (Ortega et al., 2018). Given $\{{\mathbf{v}}_{k}\}$ are orthonormal vectors, one can decompose a signal $\mathbf{x}$} into its graph frequency components as ${\bm{\alpha}}={\mathbf{V}}^{\top}{\mathbf{x}}$, where ${\mathbf{L}}^{u}={\mathbf{V}}\mathrm{diag}(\{\lambda_{k}\}){\mathbf{V}}^{\top}$, and $\mathbf{V}$$^{\top}$ is the graph Fourier transform (GFT).
The $\ell_{2}$-norm graph Laplacian regularizer (GLR) (Pang & Cheung, 2017), ${\mathbf{x}}^{\top}{\mathbf{L}}^{u}{\mathbf{x}}=\sum_{(i,j)\in{\mathcal{E}}^{u}}w^{u}_{i,j}(x_{i}-x_{j})^{2}$, is used to regularize an ill-posed signal restoration problem, like denoising, to bias low-frequency signal reconstruction (i.e., signals consistent with similarity graph ${\mathcal{G}}^{u}$) given observation $\mathbf{y}$}:
${\mathbf{x}}^{*}$ $=\arg\min_{\mathbf{x}}\|{\mathbf{y}}-{\mathbf{x}}\|^{2}_{2}+\mu\,{\mathbf{x}}^{\top}{\mathbf{L}}^{u}{\mathbf{x}}$
$=({\mathbf{I}}+\mu{\mathbf{L}}^{u})^{-1}{\mathbf{y}}={\mathbf{V}}\mathrm{diag}\left(\frac{1}{1+\mu\lambda_{1}},\ldots,\frac{1}{1+\mu\lambda_{N}}\right){\mathbf{V}}^{\top}{\mathbf{y}}.$ (1)
@@
Similar frequency notion cannot be simply defined for directed graphs via eigen-decomposition of directed graph Laplacian ${\mathbf{L}}^{d}$, because ${\mathbf{L}}^{d}$ is asymmetric, and thus the Spectral Theorem (Hawkins, 1975) does not apply. We circumvent this problem by designing new variational terms in Section 3.2.
@@ 3.2.1 $\ell_{2}$-norm Variational Term
Interpreting ${\mathbf{W}}_{r}^{d}$ as a graph shift operator (GSO) (Chen et al., 2015) on ${\mathcal{G}}^{d}$, we first define an $\ell_{2}$-norm variational term [^4] A similar directed graph variational term was defined in Li et al. (2023) towards an objective for directed graph sampling. We focus instead on directed graph signal restoration. for ${\mathcal{G}}^{d}$ as the squared difference between $\mathbf{x}$} and its graph-shifted version ${\mathbf{W}}_{r}^{d}{\mathbf{x}}$:
@@
We interpret eigen-pairs $\{(\xi_{k},{\mathbf{u}}_{k})\}$ of ${\mathcal{L}}^{d}_{r}$ as graph frequencies and graph Fourier modes for directed graph ${\mathcal{G}}^{d}$, respectively, similar to eigen-pairs $\{(\lambda_{k},{\mathbf{v}}_{k})\}$ of graph Laplacian ${\mathbf{L}}^{u}$ for an undirected graph. Importantly, we prove that our frequency definition using DGLR defaults to pure sinusoids as frequencies in the unweighted directed line graph case. Thus, our algorithm minimizing DGLR includes GSP schemes like Ramakrishna et al. (2020) that analyze signals along the time dimension using classical Fourier filters as special cases.
@@
Remark: Using our symmetrized ${\mathcal{L}}^{d}_{r}$ does not mean directionality has been lost. For the 3-node DAG in Fig. 2(c), the variational term is $\|{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|^{2}_{2}=(x_{3}-\frac{1}{2}(x_{1}+x_{2}))^{2}$ and computes to 0 for ${\mathbf{x}}=[2~~0~~1]^{\top}$. On the other hand, for the DAG in Fig. 2(d) with opposite directions, the variational term is $\|{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|^{2}_{2}=(x_{1}-x_{3})^{2}+(x_{2}-x_{3})^{2}$ and computes to 0 only if $\mathbf{x}$} is a constant vector.
@@ 3.2.2 $\ell_{1}$-Norm Variational Term
Unlike DGLR in eq. 2, DGTV is not symmetric and has no obvious frequency interpretation. Nonetheless, we provide a two-channel filterbank interpretation in the sequel.
@@ 3.3 Optimization Formulation
$\min_{\mathbf{x}}\|{\mathbf{y}}-{\mathbf{H}}{\mathbf{x}}\|^{2}_{2}+\mu_{u}{\mathbf{x}}^{\top}{\mathbf{L}}^{u}{\mathbf{x}}+\mu_{d,2}{\mathbf{x}}^{\top}{\mathcal{L}}_{r}^{d}{\mathbf{x}}+\mu_{d,1}\|{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|_{1}$ (4)
where $\mu_{u},\mu_{d,2},\mu_{d,1}\in\mathbb{R}_{+}$ are weight parameters for the three regularization terms. Combination of $\ell_{2}$- and $\ell_{1}$-norm penalties—DGLR and DGTV for directed graph ${\mathcal{G}}^{d}$ in our case [^5] We can employ both $\ell_{2}$- and $\ell_{1}$-norm regularization terms for the undirected graph ${\mathcal{G}}^{u}$ as well. For simplicity, we focus on designing new regularization terms for directed graphs in this paper. —is called elastic net regularization in statistics (Zou & Hastie, 2005) with demonstrable improved robustness. Further, employing both regularization terms means that, after algorithm unrolling, we can adapt an appropriate mixture by learning weights $\mu_{d,2}$ and $\mu_{d,1}$ per neural layer.
Convex minimization in eq. 4 is composed of smooth $\ell_{2}$-norm terms and one non-smooth $\ell_{1}$-norm term. We pursue a divide-and-conquer approach and solve eq. 4 via an ADMM framework (Boyd et al., 2011) by minimizing $\ell_{2}$- and $\ell_{1}$-norm terms in eq. 4 alternately.

### v3

@@ Abstract
Unlike conventional “black-box” transformers with classical self-attention mechanism, we build a lightweight and interpretable transformer-like neural net by unrolling a mixed-graph-based optimization algorithm to forecast traffic with spatial and temporal dimensions. We construct two graphs: an undirected graph ${\mathcal{G}}^{u}$ capturing spatial correlations across geography, and a directed graph ${\mathcal{G}}^{d}$ capturing sequential relationships over time. We predict future samples of signal $\mathbf{x}$}, assuming it is “smooth” with respect to both ${\mathcal{G}}^{u}$ and ${\mathcal{G}}^{d}$, where we design new $\ell_{2}$ and $\ell_{1}$-norm variational terms to quantify and promote signal smoothness (low-frequency reconstruction) on a directed graph. We design an iterative algorithm based on alternating direction method of multipliers (ADMM), and unroll it into a feed-forward network for data-driven parameter learning. We periodically insert graph learning modules for ${\mathcal{G}}^{u}$ and ${\mathcal{G}}^{d}$ that play the role of self-attention. Experiments show that our unrolled networks achieve competitive traffic forecast performance as state-of-the-art prediction schemes, while reducing parameter counts drastically [^1] https://github.com/SingularityUndefined/Unrolling-GSP-STForecast..
algorithm unrolling, mixed graph, graph signal processing, lightweight model, traffic forecasting
@@ 1 Introduction
Transformer, based on the classical self-attention mechanism (vaswani17attention), is now a dominant deep learning (DL) architecture that has achieved state-of-the-art (SOTA) performance across multiple fields (transformer_overview_2020; zamir2022restormer). However, like other “off-the-shelf” DL models, transformer has a massive number of parameters, and its internal structure is not easily interpretable. Model size reduction is a major industrial concern for practical computation and/or memory-constrained environments [^2] See a brief review of lightweight learning models for different applications in Appendix A.1..
Algorithm unrolling (monga21) offers an alternative hybrid model-based / data-driven paradigm; by “unrolling” iterations of a model-based algorithm minimizing an optimization objective into neural layers stacked together to form a feed-forward network for data-driven parameter learning, the resulting network can be both mathematically interpretable [^3] Common in algorithm unrolling (monga21), “interpretability” here means that each neural layer corresponds to an iteration of an optimization algorithm minimizing a mathematically-defined objective. and competitive in performance. Notably, yu23nips designs an algorithm minimizing a sparse rate reduction (SRR) objective that unrolls into a transformer-like neural net for image classification. Do2024 designs graph-based algorithms minimizing graph smoothness priors that unroll into transformer-like neural nets for image interpolation while reducing parameters. However, only a positive undirected graph model was used in Do2024 to capture simple pairwise correlations between neighboring pixels in a static image.
@@
We roughly categorize traffic forecast methods into model-based and DL-based methods. Among model-based methods, one approach to traffic forecast is to filter the signal along the spatial and temporal dimensions independently (ramakrishna2020)e.g., employ graph spectral filters in the graph signal processing (GSP) literature (ortega18ieee; cheung18) along the irregular spatial dimension, and traditional Fourier filters along the regular time dimension. Another approach is to model the spatial correlations across time as a sequence of slowly time-varying graphs, and then filter each spatial signal at time $t$ separately, where the changes in consecutive graph adjacency matrices in time are constrained by the Frobenius norm (kalofolias2017), $\ell_{1}$-norm (yamada2019), or low-rankness (bagheri2024). However, neither approach exploits spatial and temporal relationships simultaneously for optimal performance.
@@
Recent DL-based efforts address traffic forecast by building elaborate transformer architectures to capture spatial and temporal relations (feng2022; gao2022; jin2024) [^4] A detailed review of DL models on traffic forecast is included in Appendix A.2.. Like transformers used in other fields (zamir2022restormer), these are also black boxes with large parameter counts. Graph attention networks (GAT) (velickovic2018) and graph transformers (dwivedi2021) are adaptations of the self-attention mechanism in transformers for graph-structured data (e.g., computing output embeddings using only input embeddings in local graph neighborhoods). However, they are still uninterpretable and maintain large parameter sizes.
@@
In contrast, we extend Do2024 to build lightweight “white-box” transformers that are fully interpretable via mixed graph algorithm unrolling for traffic forecast. Specifically, we learn two graphs from data: i) undirected graph ${\mathcal{G}}^{u}$ to capture spatial correlations across geography, and ii) a directed graph ${\mathcal{G}}^{d}$ to capture sequential relationships over time. We show that our graph learning modules play the role of self-attention (bahdanau14), and thus our unrolled graph-based neural nets are transformers. Given the two learned graphs, we define a prediction objective for future samples in signal $\mathbf{x}$}, assuming $\mathbf{x}$} is “smooth” with respect to (w.r.t.) both ${\mathcal{G}}^{u}$ and ${\mathcal{G}}^{d}$ in variational terms: graph Laplacian regularizer (GLR) (pang17) for undirected ${\mathcal{G}}^{u}$, and newly designed directed graph Laplacian regularizer (DGLR) and directed graph total variation (DGTV) for directed ${\mathcal{G}}^{d}$. We devise a corresponding linear-time optimization algorithm based on alternating direction method of multipliers (ADMM) (boyd11) that unrolls into neural layers of a feed-forward network for parameter learning. For the first time, our proposed DGLR and DGTV variational terms quantify and promote signal smoothness on a directed graph, with intuitive low-pass filter interpretations [^5] See a detailed review of related works on directed graph Laplacians and discussion of our novelty in directed graph frequency analysis in Appendix A.3.. Extensive experiments show that our networks achieve competitive traffic forecast performance as SOTA prediction schemes, while employing drastically fewer parameters (7.2% of transformer-based PDFormer (jiang2023pdformer)).
@@ 2 Preliminaries
Denote by ${\mathcal{G}}^{u}({\mathcal{V}},{\mathcal{E}}^{u},{\mathbf{W}}^{u})$ an undirected graph with node set ${\mathcal{V}}=\{1,\ldots,N\}$ and undirected edge set ${\mathcal{E}}^{u}$, where $(i,j)\in{\mathcal{E}}^{u}$ implies that an undirected edge exists connecting nodes $i$ and $j$ with weight $w_{i,j}^{u}=W_{i,j}^{u}$, and $(i,j)\notin{\mathcal{E}}$ implies $W_{i,j}^{u}=0$. ${\mathbf{W}}^{u}\in\mathbb{R}^{N\times N}$ is the symmetric adjacency matrix for ${\mathcal{G}}^{u}$. Denote by ${\mathbf{D}}^{u}\in\mathbb{R}^{N\times N}$ a diagonal degree matrix, where $D_{i,i}^{u}=\sum_{j}W_{j,i}^{u}$. The symmetric graph Laplacian matrix is ${\mathbf{L}}^{u}\triangleq{\mathbf{D}}^{u}-{\mathbf{W}}^{u}$ (ortega18ieee). ${\mathbf{L}}^{u}$ is positive semi-definite (PSD) if all edge weights are non-negative $w^{u}_{i,j}\geq 0,\forall i,j$ (cheung18). The symmetric normalized graph Laplacian matrix is ${\mathbf{L}}^{u}_{n}\triangleq{\mathbf{I}}-({\mathbf{D}}^{u})^{-1/2}{\mathbf{W}}^{u}({\mathbf{D}}^{u})^{-1/2}$, while the asymmetric random-walk graph Laplacian matrix is ${\mathbf{L}}^{u}_{r}\triangleq{\mathbf{I}}-({\mathbf{D}}^{u})^{-1}{\mathbf{W}}^{u}$.
One can define a graph spectrum by eigen-decomposing the real and symmetric graph Laplacian ${\mathbf{L}}^{u}$ (or normalized graph Laplacian ${\mathbf{L}}_{n}^{u}$), where the $k$-th eigen-pair $(\lambda_{k},{\mathbf{v}}_{k})$ is interpreted as the $k$-th graph frequency and Fourier mode, respectively (ortega18ieee). Given $\{{\mathbf{v}}_{k}\}$ are orthonormal vectors, one can decompose a signal $\mathbf{x}$} into its graph frequency components as ${\boldsymbol{\alpha}}={\mathbf{V}}^{\top}{\mathbf{x}}$, where ${\mathbf{L}}^{u}={\mathbf{V}}\mathrm{diag}(\{\lambda_{k}\}){\mathbf{V}}^{\top}$, and $\mathbf{V}$$^{\top}$ is the graph Fourier transform (GFT).
The $\ell_{2}$-norm graph Laplacian regularizer (GLR) (pang17), ${\mathbf{x}}^{\top}{\mathbf{L}}^{u}{\mathbf{x}}=\sum_{(i,j)\in{\mathcal{E}}^{u}}w^{u}_{i,j}(x_{i}-x_{j})^{2}$, is used to regularize an ill-posed signal restoration problem, like denoising, to bias low-frequency signal reconstruction (i.e., signals consistent with similarity graph ${\mathcal{G}}^{u}$) given observation $\mathbf{y}$}:
${\mathbf{x}}^{*}$ $=\arg\min_{\mathbf{x}}\|{\mathbf{y}}-{\mathbf{x}}\|^{2}_{2}+\mu\,{\mathbf{x}}^{\top}{\mathbf{L}}^{u}{\mathbf{x}}=({\mathbf{I}}+\mu{\mathbf{L}}^{u})^{-1}{\mathbf{y}}$
$={\mathbf{V}}\mathrm{diag}\left(\frac{1}{1+\mu\lambda_{1}},\ldots,\frac{1}{1+\mu\lambda_{N}}\right){\mathbf{V}}^{\top}{\mathbf{y}}.$ (1)
@@
Similar frequency notion cannot be simply defined for directed graphs via eigen-decomposition of directed graph Laplacian ${\mathbf{L}}^{d}$, because ${\mathbf{L}}^{d}$ is asymmetric, and thus the Spectral Theorem (hawkins1975) does not apply. We circumvent this problem via new variational terms in Section 3.2.
@@ 3.2.1 $\ell_{2}$-norm Variational Term
Interpreting ${\mathbf{W}}_{r}^{d}$ as a graph shift operator (GSO) (chen15) on ${\mathcal{G}}^{d}$, we first define an $\ell_{2}$-norm variational term [^6] A similar directed graph variational term was defined in li2024 towards an objective for directed graph sampling. We focus instead on directed graph signal restoration. for ${\mathcal{G}}^{d}$ as the squared difference between $\mathbf{x}$} and its graph-shifted version ${\mathbf{W}}_{r}^{d}{\mathbf{x}}$:
@@
We interpret eigen-pairs $\{(\xi_{k},{\mathbf{u}}_{k})\}$ of ${\mathcal{L}}^{d}_{r}$ as graph frequencies and graph Fourier modes for directed graph ${\mathcal{G}}^{d}$, respectively, similar to eigen-pairs $\{(\lambda_{k},{\mathbf{v}}_{k})\}$ of graph Laplacian ${\mathbf{L}}^{u}$ for an undirected graph. Importantly, we prove that our frequency definition using DGLR defaults to pure sinusoids as frequencies in the unweighted directed line graph case. Thus, our algorithm minimizing DGLR includes GSP schemes like ramakrishna2020 that analyze signals along the time dimension using classical Fourier filters as special cases.
@@
Remark: Theorem 3.1 only gives a special case when the DGLR of a graph is equal to the GLR of its undirected base graph. In fact, using our symmetrized ${\mathcal{L}}^{d}_{r}$ does not mean directionality has been lost. As an example, for the 3-node DAG in Fig. 2(c), the variational term is $\|{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|^{2}_{2}=(x_{3}-\frac{1}{2}(x_{1}+x_{2}))^{2}$ and computes to 0 for ${\mathbf{x}}=[2~~0~~1]^{\top}$. On the other hand, for the DAG in Fig. 2(d) with opposite directions, the variational term is $\|{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|^{2}_{2}=(x_{1}-x_{3})^{2}+(x_{2}-x_{3})^{2}$ and computes to 0 only if $\mathbf{x}$} is a constant vector. With our directed graph contruction with time windows as described in Section 3.1, since all nodes after the second time step has more than one predecessors, the DGLR of the undirected graph does not equal the GLR of its base graph.
@@ 3.2.2 $\ell_{1}$-Norm Variational Term
Unlike DGLR in Equation 2, DGTV is not symmetric and has no obvious frequency interpretation. Nonetheless, we provide a two-channel filterbank interpretation in the sequel.
@@ 3.3 Optimization Formulation
$\hskip-4.25006pt\min_{\mathbf{x}}\|{\mathbf{y}}\!-\!{\mathbf{H}}{\mathbf{x}}\|^{2}_{2}+\mu_{u}{\mathbf{x}}\!^{\top}{\mathbf{L}}\!^{u}{\mathbf{x}}+\mu_{d,2}{\mathbf{x}}\!^{\top}\!{\mathcal{L}}_{r}^{d}{\mathbf{x}}+\mu_{d,1}\!\|{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|_{1},$ (4)
where $\mu_{u},\mu_{d,2},\mu_{d,1}\in\mathbb{R}_{+}$ are weight parameters for the three regularization terms. Combination of $\ell_{2}$- and $\ell_{1}$-norm penalties—DGLR and DGTV for directed graph ${\mathcal{G}}^{d}$ in our case [^7] We can employ both $\ell_{2}$- and $\ell_{1}$-norm regularization terms for the undirected graph ${\mathcal{G}}^{u}$ as well. For simplicity, we focus on designing new regularization terms for directed graphs here. —is called elastic net regularization in statistics (zou2005) with demonstrable improved robustness. Further, employing both regularization terms means that, after algorithm unrolling, we can adapt an appropriate mixture by learning weights $\mu_{d,2}$ and $\mu_{d,1}$ per neural layer.
Convex minimization in Equation 4 is composed of smooth $\ell_{2}$-norm terms and one non-smooth $\ell_{1}$-norm term. We pursue a divide-and-conquer approach and solve Equation 4 via an ADMM framework (boyd11) by minimizing $\ell_{2}$- and $\ell_{1}$-norm terms in Equation 4 alternately.